OpenStudy (anonymous):

Solve linear equations. 3/17x=12

OpenStudy (saifoo.khan):

\[\frac{3}{17}x = 12\] \[x = \frac{12 \times 17}{3}\] \[x = 4 \times 17\]

OpenStudy (anonymous):

Ok that was correct! is there a formula for this?

OpenStudy (saifoo.khan):

Nope, no particular formula!

OpenStudy (anonymous):

where did the 4 come from

OpenStudy (saifoo.khan):